# GRAND: Graph Release with Assured Node Differential Privacy

### v2

@@ 4.1 The Prototype Node DP Mechanism: An Oracle Scenario
We introduce the principle of our design regarding how it may preserve network properties. The prototypical idea is to (i) acquire a latent vector for each node, (ii) privatize the latent vectors while preserving their distribution, and (iii) generate a private network through using the private latent vectors. The private network is expected to have the same network properties as the original one, since they are generated from latent vectors of the same distribution.
@@
Under the latent space model in Definition 2, recall that the latent vectors $Z_{i}=(Z_{i1},\ldots,Z_{id})^{T}\in\mathbb{R}^{d},i\in[N]$ are an i.i.d sample from $F$. To introduce our prototype design, we assume that we know two pieces of oracle information.
@@
  • The joint cumulative distribution function (CDF) $F$ of the true latent vectors: From $F$, we can also derive all conditional distributions. Let $F^{\,l\mid 1:(l-1)}(\cdot|Z_{i1},\ldots,Z_{i,l-1})$ be the conditional CDF of $Z_{il}$ given $(Z_{i1},\ldots,Z_{i,l-1})$. When the context is clear, we write $F^{l|1:(l-1)}$ as $F^{l}$ for simplicity. In particular, let the marginal CDF of the first coordinate be $F^{1}$.
@@
  • The true latent vectors $Z_{i},i\in[N]$. Furthermore, let $\mathbf{Z}_{l}=(Z_{1l},\ldots,Z_{Nl})^{T}$ be the $l$th latent vector of all nodes for $l\in[d]$. For each $l$, we can treat $\mathbf{Z}_{l}$ as a sequence of univariate observations.
@@
We next introduce a distribution-invariant privacy mechanism (DIP) proposed in Bi and Shen (2023) to perturb the univariate data to achieve differential privacy, which will then be applied to all coordinates by conditioning. Specifically, applying $F^{1}$ to $\mathbf{Z}_{1}$, we note that $F^{1}(Z_{i1})$ follows Uniform(0,1). We then perturb $F^{1}(Z_{i1})$ by adding an independent noise $e_{i1}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\text{Laplace}(0,1/\varepsilon)$ for $i\in[N]$. Following the standard Laplace mechanism, we know that $F^{1}(Z_{i1})+e_{i1}$ satisfies $\varepsilon$-differential privacy (Dwork, 2006). Let $G$ be the CDF of $F^{1}(Z_{i1})+e_{i1}$ whose expression can be exactly obtained via convolution. Applying $G$ to $F^{1}(Z_{i1})+e_{i1}$ also results in a uniform random variable. Then the privatization mechanism $\mathfrak{m}^{1}$ for $\mathbf{Z}_{1}$ can be expressed as
$\tilde{Z}_{i1}\equiv\mathfrak{m}_{1}(Z_{i1})=(F^{1})^{-1}\circ G\big{(}F^{1}(Z_{i1})+e_{i1}\big{)},\quad i=1,\ldots,N,$
@@
where $\circ$ denotes function composition. We can see that $\tilde{Z}_{i1}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}F^{1}$. Meanwhile, $\tilde{Z}_{i1}$ is differentially private (Dwork, 2006).
@@
For $l\geq 2$, we apply the same strategy to privatize each latent vector sequentially, using probability chain rule (Schum, 2001). Let $F^{l}(\cdot):=F^{l}(\cdot|Z_{i1},\ldots,Z_{i,l-1})$ and $\tilde{F}^{l}(\cdot):=F^{l}(\cdot|\tilde{Z}_{i1},\ldots,\tilde{Z}_{i,l-1})$. Then
$\tilde{Z}_{il}\equiv\mathfrak{m}_{l}(Z_{il}|\tilde{Z}_{i1},\ldots,\tilde{Z}_{i,l-1})=(\tilde{F}^{l})^{-1}\circ G\big{(}F_{l}(Z_{il})+e_{il}\big{)},$ (3)
@@
where $\mathfrak{m}_{l}(\cdot)$ denotes the privatization process for $\mathbf{Z}_{l}$, $l=2,\ldots,d$. Notice that, since $Z_{i}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}F$, we have $F^{l}(Z_{il})\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\text{Uniform(0,1)}$. On the other hand, the use of $(\tilde{F}^{l})^{-1}$ in (3) is to guarantee differential privacy of the preceding variables and preserve that $\tilde{Z}_{i}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}F$. For notational simplicity, we write
@@
where $\tilde{Z}_{i}=(\tilde{Z}_{i1},\ldots,\tilde{Z}_{id})^{T}$, $\mathbf{e}_{i}=(e_{i1},\ldots,e_{id})^{T}$, and $\mathfrak{m}()$ denotes the sequential application of $\mathfrak{m}_{1},\ldots,\mathfrak{m}_{d}$. Here $\tilde{Z}_{i}$ is the differentially private perturbation of $Z_{i}$.
Subsequently, we can follow Definition 2 and use the same generative function $W$ to generate the private adjacency matrix $\tilde{A}$. Then given the fact that $\tilde{Z}_{i}$ follows $F$ and that $W$ is known, $\tilde{A}$ should follow the same distribution as $A$. This ensures that network properties of $\tilde{A}$, especially those that can be presented as summary statistics, remains the same. Meanwhile, due to the fact that $\tilde{Z}_{i}$ satisfies $\varepsilon$-differential privacy, $\tilde{A}$ follows NDP.
@@ 4.2 The Proposed Network Release with Node-wise Estimation
    In our prototype procedure (e.g., see (3)), the outermost layer of operation depends on the inverse CDF. This is the final layer of operation without additional privacy protection. If the estimated CDF is based on the nodes to be released, the privacy of these nodes cannot be protected.
@@
    In practice, fitting any network models almost always requires additional tuning or model selection. Rigorously speaking, once such procedure involves the to-be-released data, DP cannot be guaranteed unless a specialized tuning procedure is designed.
To rectify these challenges, we assume that we have a network $A$ between $N=m+n$ nodes available. We would like to hold out the network of $m$ nodes while privatizing and releasing only the network between the remaining $n$ nodes. By doing this, we will be able to use the hold-out network as the reference to obtain estimates about $F$ and $Z_{i}$’s and also for tuning if needed while ensuring the node DP for the released network between the $n$ nodes. Figure 1 gives a high-level flowchart demonstrating the proposed method, whose details will be introduced soon. Our data splitting practice is also seen in the graph neural network literature (e.g., Yang et al., 2016; Kipf and Welling, 2017), where a sub-network is used for model training, and another sub-network is used for testing/evaluation. In statistical literature, such a data splitting strategy is also used by Chen and Lei (2018) for community detection problems.
@@
Figure 1: The proposed privacy-preserving scheme illustrated: An $N\times N$ adjacency matrix AA is used as the input, and a private version of a $n\times n$ subnetwork $A^{11}$, namely $\tilde{A}^{11}$, is generated as the output. All privacy vulnerable quantities are colored in red. Each of the latent vectors of nodes $i\leq n$ is separately estimated by a node-wise estimation with the help of hold-out estimates $\{\hat{Z}_{i}\}_{i>n}$. Then $\{\hat{Z}_{i}\}_{i\leq n}$ goes through the DIP mechanism to encode differential privacy. Subsequently, the privatized subnetwork $\tilde{A}^{11}$ is generated from the private latent vectors $\{\tilde{Z}_{i}\}_{i\leq n}$.
Specifically, assuming the network is generated from a general latent space model, we apply one of the standard model estimation procedures discussed in Remark 1 to $A^{22}$. This step would result in an estimate of the latent positions for the hold-out nodes, denoted by $\{\hat{Z}_{i}\}_{i=n+1}^{n+m}$. From this step, we can estimate the corresponding CDF $F$ for the latent vectors $\{\hat{Z}_{i}\}_{i=n+1}^{n+m}$. This can be achieved through following standard practices such as empirical CDFs, kernel smoothing, splines, or $K$-nearest neighbors (Wasserman, 2006). Denote this estimated $F$ by $\hat{F}$.
@@
For most commonly used $W$, this problem is easy to solve using standard algorithms. In many particular cases, the problem can be significantly simplified. For example, under the inner-product latent space model (Example 1), (5) becomes a simple logistic regression problem:
$\hat{Z}_{i}=\mathop{\arg\min}_{(X_{i},\alpha_{i})\in{\bf R}^{d}}\sum_{j=n+1}^{n+m}\big{\{}-A_{ij}\log\sigma(\alpha_{i}+\hat{\alpha}_{j}+X_{i}^{T}\hat{X}_{j})-(1-A_{ij})\log(1-\sigma(\alpha_{i}+\hat{\alpha}_{j}+X_{i}^{T}\hat{X}_{j}))\big{\}}.$ (6)
Other reasonable objectives can be used as alternatives to the negative log-likelihood in (5), and these alternatives may further simplify the algorithm. For example, under the RDPG (Example 2), it is easy to use the sum of squared errors for moment matching as the M-estimation criterion rather than the negative log-likelihood. This is because the estimation becomes a node-wise linear regression procedure.
$\hat{Z}_{i}=\arg\min_{Z_{i}\in{\bf R}^{d}}\sum_{j=n+1}^{n+m}(A_{ij}-Z_{i}^{T}\hat{Z}_{j})^{2}.$ (7)
With the estimated $\hat{F}$ and $\{\hat{Z}_{i}\}_{i\in[n]}$, we can apply the prototypical procedure in Section 4.1 for privatization. In particular, $\hat{F}$ does not utilize any information from the to-be-released nodes, thus can be used to replace $F$. Calling the formula in (4), for each $i\in[n]$, we have
@@
Subsequently, we generate the private network $\tilde{A}^{11}$ based on $\{\tilde{Z}_{i}\}_{i\in[n]}$ following the same general latent space model. Algorithm 1 provides a summary of the proposed method.
Algorithm 1 The proposed GRAND mechanism by node-wise estimation
@@
3: Estimate the hold-out latent vectors $\{\hat{Z}_{i}\}_{i=n+1}^{n+m}$ based on $A^{22}$ using a standard network model estimation procedure as provided in Remark 1.
@@
Similar to the requirements on the holdout dataset in DIP (Bi and Shen, 2023), the holdout set should not be accessed, altered, queried, or released, and should be deleted immediately once Algorithm 1 is implemented. Alternatively, additional approaches beyond the discussion provided in Bi and Shen (2023) may also be viable. First, we note that $\hat{F}$ is the estimated CDF of the latent vectors. With additional regularity conditions, one can apply some particular mechanisms, including the Private Signed Measure Mechanism (He et al., 2023), which generate $\hat{F}$ that are both differentially private and asymptotically consistent.
Second, it is also possible in the social network context that different nodes have different preferences in privacy protection. In such a case, one can construct the holdout set using nodes that are less sensitive to data release. On the one hand, one can apply differential privacy for functions (Hall et al., 2013) and achieves $(\varepsilon,\delta)$-differential privacy for a kernel density built on the holdout set. On the other hand, to accommodate the possible distributional discrepancy between the holdout and the release sets, a transfer learning model as discussed in Pan and Yang (2010) and Zhuang et al. (2020) can be used to replace (5). Nevertheless, the adoption of both a $(\varepsilon,\delta)$-differentially private holdout set and a transfer learning model entails additional theoretical establishment, which is acknowledged in future work in Section 8.
@@
In the estimation procedure, it can be seen that $A^{11}$ is not used. This is to avoid the privacy spill-over issue. Therefore, from a statistical estimation perspective, the estimation is essentially based on a size-$(m+1)$ network rather than a size-$N$ network. which will be explicitly verified in our theory and numerical results. The statistical efficiency reduction from size $N$ to size $m$ is the price we pay for the node-level privacy.
Remark 6.
The computational bottleneck of the algorithm is on the standard model estimation procedure in Step 3, which depends on $m$. The node-wise estimation is typically trivial. Moreover, it can be done completely in parallel, which further improves the computational feasibility.
Remark 7 (Selection of the private network size).
Intuitively, we hope to maximize the released size $n$, subject to m=Nnm=N-n being large enough for model estimation. In practice, this choice may also depend on other factors, such as the target size of the released networks, computational resources and differences in privacy levels for subjects in the network. In our numerical studies, we use $n=0.5N$ for simplicity.
Remark 8.
While we focus on undirected networks in this paper, it is not difficult to see that the procedure is ready to be extended to directed networks. In the directed networks, the corresponding latent space model would associate with each node a “sending” latent vector and a “receiving” latent vector. We will have to estimate both latent vectors for each node, using both $A^{12}$ and $A^{21}$, and then combine the “sending” and “receiving” vectors together in the privatization step.

### v3

@@ 4.1 The Prototype Node DP Mechanism: An Oracle Scenario
We introduce the principle of our design regarding how it may preserve network properties. The prototypical idea is to (i) acquire a latent vector for each node, (ii) privatize the latent vectors while preserving their distribution, and (iii) generate a private network through using the private latent vectors. The private network is expected to have the same network properties as the original one, since share the same latent vector distribution.
@@
Under the latent space model in Definition 2, recall that the latent vectors satisfy $Z_{i}=(Z_{i1},\ldots,Z_{id})^{\top}\in\mathbb{R}^{d}$, $i\in[N]$ and are an i.i.d. sample from $F$. To introduce our prototype design, we assume that we know two pieces of oracle information.
@@
  • One is the joint cumulative distribution function (CDF) $F$ of the true latent vectors. From $F$, we can also derive all conditional distributions. Let $F^{l\,|\,1:(l-1)}(\cdot\,|\,Z_{i1},\ldots,Z_{i,l-1})$ be the conditional CDF of $Z_{il}$ given $(Z_{i1},\ldots,Z_{i,l-1})$. When the context is clear, we write $F^{l\,|\,1:(l-1)}$ as $F^{l}$ for simplicity. In particular, let the marginal CDF of the first coordinate be $F^{1}$.
@@
  • The other is the true latent vectors $Z_{i}$, $i\in[N]$. Furthermore, let $\mathbf{Z}_{l}=(Z_{1l},\ldots,Z_{Nl})^{\top}$ be the $l$-th latent vector of all nodes for $l\in[d]$. For each $l$, we can treat $\mathbf{Z}_{l}$ as a sequence of univariate observations.
@@
We next introduce a distribution-invariant privacy mechanism (DIP) proposed in Bi and Shen (2023) to perturb the univariate data to achieve differential privacy, which will then be applied to all coordinates by conditioning. Specifically, applying $F^{1}$ to $\mathbf{Z}_{1}$, we note that $F^{1}(Z_{i1})$ follows $\text{Uniform}(0,1)$. We then perturb $F^{1}(Z_{i1})$ by adding an independent noise $e_{i1}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\text{Laplace}(0,1/\varepsilon)$ for $i\in[N]$. Following the standard Laplace mechanism, we know that $F^{1}(Z_{i1})+e_{i1}$ satisfies $\varepsilon$-differential privacy (Dwork, 2006). Let $G$ be the CDF of $F^{1}(Z_{i1})+e_{i1}$ whose expression can be exactly obtained via convolution. Applying $G$ to $F^{1}(Z_{i1})+e_{i1}$ also results in a uniform random variable. The privatization mechanism $\mathfrak{m}^{1}$ for $\mathbf{Z}_{1}$ can be written as
$\tilde{Z}_{i1}\equiv\mathfrak{m}_{1}(Z_{i1})=(F^{1})^{-1}\circ G(F^{1}(Z_{i1})+e_{i1}),\quad i=1,\ldots,N,$
@@
where $\circ$ denotes function composition. We can see that $\tilde{Z}_{i1}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}F^{1}$. Meanwhile, $\tilde{Z}_{i1}$ is differentially private (Dwork, 2006).
@@
For $l\geq 2$, we apply the same strategy to privatize each latent vector sequentially, using probability chain rule (Schum, 2001). Let $F^{l}(\cdot):=F^{l}(\cdot\,|\,Z_{i1},\ldots,Z_{i,l-1})$ and $\tilde{F}^{l}(\cdot):=F^{l}(\cdot\,|\,\tilde{Z}_{i1},\ldots,\tilde{Z}_{i,l-1})$.
$\tilde{Z}_{il}\equiv\mathfrak{m}_{l}(Z_{il}\,|\,\tilde{Z}_{i1},\ldots,\tilde{Z}_{i,l-1})=(\tilde{F}^{l})^{-1}\circ G(F^{l}(Z_{il})+e_{il}),$ (3)
@@
where $\mathfrak{m}_{l}(\cdot)$ denotes the privatization process for $\mathbf{Z}_{l}$, $l=2,\ldots,d$. Notice that, since $Z_{i}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}F$, we have $F^{l}(Z_{il})\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\text{Uniform}(0,1)$. On the other hand, the use of $(\tilde{F}^{l})^{-1}$ in (3) is to guarantee differential privacy of the preceding variables and preserve that $\tilde{Z}_{i}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}F$. For notational simplicity, we write
@@
where $\tilde{Z}_{i}=(\tilde{Z}_{i1},\ldots,\tilde{Z}_{id})^{\top}$, $\mathbf{e}_{i}=(e_{i1},\ldots,e_{id})^{\top}$, and $\mathfrak{m}(\cdot)$ denotes the sequential application of $\mathfrak{m}_{1},\ldots,\mathfrak{m}_{d}$. Here $\tilde{Z}_{i}$ is the differentially private perturbation of $Z_{i}$.
Subsequently, we can follow Definition 2 and use the same generative function $W$ to generate the private adjacency matrix $\tilde{A}$. Then given the fact that $\tilde{Z}_{i}$ follows $F$ and that $W$ is known, $\tilde{A}$ should follow the same distribution as $A$. This ensures that network properties of $\tilde{A}$, especially those that can be presented as summary statistics, remains the same. Meanwhile, as $\tilde{Z}_{i}$ satisfies $\varepsilon$-differential privacy, $\tilde{A}$ follows NDP.
@@ 4.2 The Proposed Network Release with Node-wise Estimation
    In our prototype procedure (e.g., see (3)), the outermost layer of operation depends on the inverse CDF. This is the final layer of operation without additional privacy protection. If the estimated CDF is based on the nodes to be released, the differential privacy of these nodes may not be guaranteed..
@@
    In practice, fitting any network models almost always requires additional tuning or model selection. Rigorously speaking, once such procedure involves the to-be-released data, DP cannot be guaranteed.
To rectify these challenges, we assume that we have a network $A$ between $N=n+m$ nodes available. We would like to hold out the network of $m$ nodes while privatizing and releasing only the network between the remaining $n$ nodes. By doing this, we will be able to use the hold-out network as the reference to obtain estimates about $F$ and $Z_{i}$’s and also for tuning if needed while ensuring the node DP for the released network between the $n$ nodes. Figure 1 gives a high-level flowchart demonstrating the proposed method, whose details will be introduced soon. Our data splitting practice is also seen in the graph neural network literature (e.g., Kipf and Welling, 2017), where a sub-network is used for model training, and another sub-network is used for testing. In statistics, the splitting strategy is also used by Chen and Lei (2018).
@@
Assuming the network is generated from a general latent space model, we apply one of the standard model estimation procedures based on the specific assumed model to $A^{22}$. For example, if an inner product latent space model is assumed, one can apply the likelihood-based gradient descent algorithm of Ma et al. (2020); Li et al. (2023a). If the RDPG is assumed, then the adjacency spectral embedding (ASE) of Sussman et al. (2014) can be used. This step would result in an estimate of the latent positions for the hold-out nodes, denoted by $\{\hat{Z}_{i}\}_{i=n+1}^{n+m}$. From this step, we can estimate the corresponding CDF $F$ for the latent vectors $\{\hat{Z}_{i}\}_{i=n+1}^{n+m}$. This can be achieved through following standard practices such as empirical CDFs, kernel smoothing, splines, or $K$-nearest neighbors (Wasserman, 2006). Denote this estimated $F$ by $\hat{F}$.
Figure 1: The proposed privacy-preserving scheme illustrated: An $N\times N$ adjacency matrix AA is used as the input, and a private version of a $n\times n$ subnetwork $A^{11}$, namely $\tilde{A}^{11}$, is generated as the output. All privacy vulnerable quantities are colored in red. Each of the latent vectors of nodes $i\leq n$ is separately estimated by a node-wise estimation with the help of hold-out estimates $\{\hat{Z}_{i}\}_{i>n}$.
@@
For most commonly used $W$, this problem is easy to solve using standard algorithms. In many particular cases, the problem can be significantly simplified. For example, under the inner product latent space model, (5) becomes a simple logistic regression problem:
$\mathop{\arg\min}_{(X_{i},\alpha_{i})\in\mathbb{R}^{d}}\sum_{j=n+1}^{n+m}\{-A_{ij}\log\sigma(X_{i}^{\top}\hat{X}_{j}+\alpha_{i}+\hat{\alpha}_{j})-(1-A_{ij})\log(1-\sigma(X_{i}^{\top}\hat{X}_{j}+\alpha_{i}+\hat{\alpha}_{j}))\}.$ (6)
Other reasonable objectives can be used as alternatives to the negative log-likelihood in (5), and these alternatives may further simplify the algorithm. For example, under the RDPG, it is easier to use the sum of squared errors as an M-estimation criterion rather than the likelihood. Thus the estimation becomes a node-wise OLS linear regression $\mathop{\arg\min}_{Z_{i}\in\mathbb{R}^{d}}\sum_{j=n+1}^{n+m}(A_{ij}-Z_{i}^{\top}\hat{Z}_{j})^{2}$.
With the estimated $\hat{F}$ and $\{\hat{Z}_{i}\}_{i\in[n]}$, we can apply the prototypical procedure in Section 4.1. In particular, $\hat{F}$ does not utilize any information from the to-be-released nodes, thus can be used to replace $F$. Using (4), for each $i\in[n]$, we have
@@
Subsequently, we generate the private network $\tilde{A}^{11}$ based on $\{\tilde{Z}_{i}\}_{i\in[n]}$ following the Bernoulli. Algorithm 1 provides a summary of the proposed method.
Note that if one has node-level attributes $U_{i}$ in addition to the network, we can simply let $\hat{Z}_{i,\rm new}=(\hat{Z}_{i},U_{i})$ and equation (7) can directly be applicable to $\hat{Z}_{i,\rm new}$. Consequently, our method and the theoretical results can be trivially extended to handle networks with node attributes. We briefly outline this extension in Appendix G.
Algorithm 1 The GRAND mechanism by node-wise estimation
@@
3: Estimate the hold-out latent vectors $\{\hat{Z}_{i}\}_{i=n+1}^{n+m}$ based on $A^{22}$ using a standard network model estimation procedure.
@@
Remark 3 (Privacy of the Holdout Set).
Similar to the requirements discussed in Bi and Shen (2023), the holdout set should not be accessed, altered, queried, or released, and should be deleted immediately once Algorithm 1 is implemented. More details can be found in Bi and Shen (2023).
We acknowledge that the holdout set does not receive the same formal node DP guarantee as the released set. However, we argue that this concession is necessary to resolve a fundamental conflict inherent to network data: the privacy spill-over effect. Intuitively, preserving network structures (e.g, by model fitting) requires exploiting pairwise relations. Conversely, node differential privacy mandates that every single node should be separable in privatization operations. These two objectives are intrinsically contradictory. Consequently, previous attempts in literature to achieve node DP on the entire graph have resulted in either a prohibitive loss of utility (network structures) or computational intractability.
Our method reconciles this tension by using the holdout set as a structural anchor. The proposed node-wise estimation (Equation (5)) decouples the released nodes in model fitting, thereby asymptotically breaking the trade-off between privacy and utility for the released network. We feel that the introduction of such a holdout set is a reasonable and necessary price to pay for the first computationally feasible mechanism that delivers both assured Node Differential Privacy and structural consistency. In Section 8, we discuss possible future directions to further improve the privacy protection for the holdout set.
Remark 4 (Tradeoff for the Private Network Size).
It can be seen that $A^{11}$ is not used in estimation. This is to avoid the privacy spill-over issue. Therefore, from a statistical estimation perspective, the estimation is essentially based on a size-$(m+1)$ network rather than a size-$N$ network. which will be explicitly verified in our theory and numerical results. The statistical efficiency reduction from size $N$ to size $m$ is the price we pay for the node-level privacy. Meanwhile, the computational bottleneck of the algorithm is on the standard model estimation procedure in Step 3 based on $m$. The node-wise estimation is typically trivial and can be completely in parallel. Intuitively, we hope to maximize the released size $n$, subject to m=Nnm=N-n being large enough for model estimation. In practice, this choice may also depend on other factors, such as the target size of the released networks and computational resources. In our numerical studies, we use $n=0.5N$ for simplicity.
@@
It is not difficult to extend the procedure for directed networks. In the directed case, the latent space model would associate with each node a “sending” latent vector and a “receiving” latent vector. We will have to estimate both latent vectors for each node, using both $A^{12}$ and $A^{21}$. The rest operations would remain similar.